OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). Help~

OpenStudy (anonymous):

First you need to find the slope of the original line. In this case it is -3

OpenStudy (anonymous):

Then you need to find the b for the equation of the second line (y=mx+b)

OpenStudy (anonymous):

sub in your x and y from the points (x=-1 y=6) and the slope from the original line (since you want them parallel) so you get y=mx+b 6=(-3)(-1)+b b=3 Then write your new equation which should be y= -3x +3